OpenStudy (anonymous):

Solve square root of x plus 14 end root minus 12 equals x

OpenStudy (anonymous):

add 12 to both sides first \[\sqrt{x+14}=x+12\]

OpenStudy (anonymous):

And then square both sides to eliminate the square root x+14=(x+12)^2

OpenStudy (anonymous):

x = –10 and x = –13 ?

OpenStudy (anonymous):

Yes